Josef owns a 35 inch TV (that is, it measures 35 inches on the diagonal). If the television is 28 inches across the bottom, find the height of the TV. (Remember a TV screen is a rectangle)
Answer by anteater(7)

Use the Pythagorean Theorem: a^2 + b^2 = c^2
In this case, the lengths of the hypotenuse (35") and one of the legs (28") are given. Plug the information into the formula, and solve to find the length of the remaining leg:
28^2 + b^2 = 35^2
784 + b^2 = 1225
Subtract 784 from both sides of the equation to isolate b^2:
b^2 = 1225 - 784
b^2 = 441 square inches
Take the square root of both sides of the equation to find b:
b = 21 inches